OpenStudy (surryyy):

If you toss 2 dice 10,000 times, how many times would you expect that " he sum of the faces equals 4"\ happens? Answer 733 550 333 833

OpenStudy (queelius):

First, how many ways are there for two die to sum to four? 1 3 2 2 3 1 So, 3 out of 6*6=36 of the possible die rolls equals four. That is, P[sum of 4] = 3/36 = 1/12 So, now, we only need to calculate P[sum of 4] * 10,000.

OpenStudy (surryyy):

so the final answer would be 833?